Text Size:
Smaller Text Normal Text Larger Text

Facebook page

Twitter profile

YouTube channel

NOTCam News (Status updates)

2023-09-06 Pumping while cold NOTCam was put on the pump while being kept cold to improve the vacuum.
2021-11-03 NOTCam Opening
03 - 29 Nov 2021
The vacuum was bad and it was decided to warm up and open. NOTCam was transported to the SB and a leak test was done before opening, but with no clear results. No apparent cause of the vacuum loss was found inside after opening and inspection. The charcoal was removed and baked, and o-rings were cleaned. Waiting as long as possible for a CaF2-sandwiched KDP to arrive (for pre-NTE testing), but eventually NOTCam needed to be closed and was set to pumping.
2021-07-16 New KDP tests The KDP crystals were remounted in NOTCam in order to test again their behaviour during cooling down. Check the report here: New KDP test report.
2021-06-22 NOTCam Opening
22 Jun - 14 Jul 2021
This was a scheduled opening of NOTCam in order to unmount 7 filters and pack them for transport to the IAC for transmission measurements as requested by the NTE team and to re-install two KDPs for NTE tests. The filters when back from the IAC were re-installed in the same wheel slots as before. The spare H and K filters that had been in the stop-wheel were dismounted and sent to Copenhagen for NTE use. The aperture wheel was checked, the pressure sensors were cleaned, and the charcoal getter was baked.
Two KDPs arriving from Copenhagen were installed in the aperture wheel slot 2 and 9 in an environment in the clean room of 30 degrees Celcius and a relative humidity of 14%. The time from opening the KDP boxes to having NOTCam closed and pumping was limited to < 3 hours. The idea was to redo the KDP tests from 2016, but now with a long warm pumping before cooling, in order to avoid any condensations on the KDPs, the hypothesis.
2020-11-11 New trolley wheels The NOTCam trolley was upgraded with new wheels, bigger and harder, so now the heavy NOTCam instrument can roll over the dome floor again.
2020-08-19 New instrument computer NOTCam was upgraded with a new computer "tessa" as a replacement of the previous instrument computer "marissa". Note that the observing system is started from "selena". The Cookbook has been updated.
2020-04-16 Warm-up on pump NOTCam was set to warm-up while being connected to the vacuum pump on April 16th. The pumping was going on until the temperature inside NOTCam was above zero, and that took 20 days. Then the LN2 tank was pumped with the 2nd turbo pump, using silica gel to avoid water in the pump. After this the tank was flushed with dry air. The PTR was started again on May 4th, and LN2 cooling started on May 5th. The vacuum pumping was continuing all the time until NOTCam reached operating temperature. This was found to be crucial, as the LN2 filling in a warm cryostate usually caused vacuum losses due to the design of the fill/vent tubes. The pressure inside NOTCam is now typically 4E-05 mbar and the holding time is 6 days with the cold table at a stable temperature until the tank gets emptied of LN2, i.e. a substantial improvement. See TMS plot May 6th - Jun 20th compared to the situation earlier this year: TMS plot Feb 6th - Mar 20th where the pressure was typically 2E-04 mbar and the cold table temperature was never stable.
2020-01-12 Less frequent LN2 fillings The limiting temperature for re-filling with LN2 was relaxed to -192 deg allowing for less frequent fillings and better behaviour of the cryostate, which can be seen in this plot covering before and after the change: TMS plot Dec 10 - Feb 05. The history behind the frequent fillings was that of a maximum precaution against a potential inaccessability of the observatory, to avoid NOTCam warming up. Some years ago we had set a strict limit (-196 deg) for the detector temperature for when to alert staff to fill NOTCam. It was, however, over years realized that the LN2 tank absorbs water, which forms ice, causing problems with LN2 filling, only small mounts could be fitted in the tank, and eventually needing a warm-up and opening. In Jan 2020 we decided to relax on the refillings and let it warm up much more before a new filling was made. This improved the cryostate behaviour.
2019-07-22 Warm-up on pump NOTCam was set to warm-up while being connected to the vacuum pump on July 22nd. The aim was to improve the holding time by pumping the LN2 tank which is suspected to have water ice accumulated in it. This is a similar procedure as the one done in May 2013 for the first time. See Cold table TMS plot. The PTR was started again on Aug 7th and LN2 filling started Aug 9th. Due to outgassing the pressure sensor got saturated and stopped showing reliable results. The pressure measured on the pump before filling with LN2 was 5.2E-6 mbar. At the next NOTCam opening the pressure sensor will be cleaned.
2018-10-30 Vacuum improvement Monitoring of the pressure inside NOTCam shows that the stability of the vacuum improved substantially after the maintenance performed during the last opening of NOTCam in June 2018. See TMS plot May-Oct 2018 of the pressure and the cold table temperature.
2018-06-11 NOTCam Opening
11 - 26 June 2018
NOTCam was set to warm-up on the morning of June 11th after observations, transported to the SB the same day, and was sufficiently warm to be opened on the 15th. The DKDP was dismounted and stored in a dry environment. The getter was set to baking. The interior was investigated for dust and wear. The LN2 tank was pumped with the roughing pump to suck out accumulated water. Pressure sensor cleaned. NOTCam was closed on the 18th and set to pump. Transport to the telescope on the 20th, PTR switched on 23rd and filled with LN2 on the 25th. Total time from warm-up to cold again estimated to about 14 days.
2018-02-19 New DKDP test
The DKDP currently mounted in NOTCam still shows the same structure as when first mounted two years ago. This means that it is either not due to ice condensation, or that water does not easily leave the KDP.
2017-05-04 PTR oil adsorber
As part of the Iwatani PTR maintenance, the oil adsorber for the compressor (model CW303) should be exchanged every 15000 running hours. From experience we have run safely about 30000 hours earlier. At 52000 hours we now contacted Kelvin International in order to refurbish this piece. The new one arrived and was installed on June 12th. The old one sent back to KIC.
2017-04-04 Spectroscopy scripts
Dithering along the slit in spectroscopic mode has been upgraded from the standard ABBA mode of 4 spectra to the new AB3 mode , which is taking 6 spectra in the sequence ABABAB. This mode allows for better background subtraction and allows one to discard the first exposure (which is useless due to the reset anomaly) with less total loss. For extended object spectroscopy the script notcam.ab3-beamswitch can be used. It is similar to AB3, but starts with an offset to the B position (such that the off-target spectrum is the one to discard) and allows for small step dithering to filter out bad pixels. For an overview of the spectroscopy scripts, see NOTCam spectroscopy scripts.
2017-03-02 OB generator
The Observing Block Generator is now offered for NOTCam, with the OB compiler fully functional for the imaging mode. From a user's point of view, the OB generator is also offered for spectroscopy. Because the compiler is still pending for this mode, the observer will run a set of template scripts in a semi-automatic mode.
2017-02-02 FITS header updates
The new NOTCam FITS header keyword IMAGECAT is introduced and set in various NOTCam template observing scripts (resetting to default value SCIENCE upon ending). The definition of FITS header keywords follow the Description of FITS headers at the NOT (2014) (PDF document).
2016-09-30 Better vacuum The cryostat vacuum is a factor of two better than what we have had over the last years. We believe this is due to the teflon-jacketed o-rings installed in the fill/vent tubes in March-2016. There seem to be no more problems with vacuum losses every time we fill with LN2.
2016-05-22 New fiber needed.
The optical fiber that goes from the detector controller sequencer board, mounted on the instrument, to the instrument computer in the electronics room was found to cause stress on the sequencer board connectors, and a replacement fiber has been ordered.
2016-04-22 Zero pixels on all quadrants.
A sudden and fatal error while observing. From one image to the next all pixels became zero valued. Trouble shooting the controller and exchanging boards fixed the problem. With this also the elevated readnoise on the upper two quadrants seemed to come down to nominal values.
2016-03-18 New LN2 filling system.
A completely new LN2 filling procedure was setup in order to avoid poking with the filling hose inside the fill tube and making damage to the interal tube. The tube seemed to have become loose on the inside. Now the filling hose from the LN2 bottle is attached with a clamp on the outside. Apart from saving the inner tube from more damage, the new system seems to improve the problem with dripping.
2016-03-15 Slow cooling down NOTCam was cooled down all the way to -180 degrees only by the aid of the PTR. This gradual cool-down took a much longer time, but was made to see if the KDPs acted differently, and also to find out at what temperature the KDP crystal would become structured. It was also expected that having NOTCam already cooled, could cause less harm to the fill/vent o-rings when finally filling the LN2, but this made no difference.
2016-03-10 The funny KDPs!
Two KDPs were tested in NOTCam cold. The crystals looked broken when cold, but both regained their transparent and smooth surface when warmed up again. Internal KDP test report.
2016-03-08 NOTCam Opening
25 Feb - 22 Mar 2016
NOTCam was set to warm up on February 25th. New O-rings for the LN2 tank fill/vent tubes had been ordered and arrived just in time for installation. NOTCam was then transported to the service building clean room on March 7th and opened. The KDPs were dismounted and put in a dry environment. The getter was set to baking. Filter-wheel 1 was examined and a not-well-solded motor cable was identified and fixed. One KDP was put back for further testing inside NOTCam. NOTCam was closed and set to pumping March 8th. Transported to the dome on March 14th, PTR switched on March 15th for a slow cool-down, and filled with LN2 on March 21st.
2016-01-15 NOTCam Opening
23 Dec - 22 Jan 2016
NOTCam was set to warm up on December 23rd. Transported to the service building clean room on January 11th. The pressure sensor was cleaned, the getter set to baking, and all gear-wheel bearings and flexible couplings checked carefully. On January 15th the spare J-band filter, mounted in the stop wheel, was taken out to be sent for a proper leak measurement. Two test blockers, one KDP and one DKDP, were installed in the small slots in the aperture wheel. They were exposed to a relative humidity of about 32% in the clean-room for about 1 hour before NOTCam was closed and set to pumping on Jan 15th, transported to the dome on Jan 18th, PTR on Jan 20th and ready cold on Jan 22nd.
2016-01-15 The KDPs arrived!
In the very last moment before closing NOTCam, the KDPs arrived to La Palma.
2015-11-01 Two years cold!
NOTCam has been two years continuously running without an opening. This is a new record. The vacuum starts to deteriorate slightly, however, so we plan for an opening relatively soon to bake the charcoal getter and vacuum pump it warm.
2015-06-01 One quadrant failed
The upper left quadrant was not responding during the observations on June 1st. Fixed before the following NOTCam run.
2015-05-29 New Argon lamp power supply installed
The new Oriel 6046 power supply for the Argon lamp arrived and was installed and tested.
2015-02-04 New PTR hose installed
The 2nd new PTR hose of ultra-flex type arrived on December 12th and was connected to NOTCam on February 4th. From now on we are running with two hoses of the same ultra-flex type. The old hoses are spares.
2015-01-28 Oriol Argon lamp power supply ordered
It was realized that the reason for the Argon lamp switch-on delay could be in the power supply we use. Thanks to kind help from the TNG, it was found that our Argon lamp works well with their power supply. Therefore we decided to order a new power supply of model Oriol 6046 (plus spare).
2014-12-03 IRAF package upgraded to v2.6
The (quicklook) reduction package for NOTCam imaging has been upgraded to v2.6. This mainly involves including rejection masks to better handle the noise properties in the combined image where the number of overlaps differ, plus some bug fixes. Download and installation info.
2014-11-07 Z and Y band spectroscopy in ETC
The ETC was updated (v2.4) to include NOTCam spectroscopy in the Z (0.830-0.927 micron) and Y (0.978-1.081 micron) bands . The numbers are based on measured throughput from 2011 using the star SP2209+178 and the 1mm large pinhole.
2014-09-29 New script: notcam.acquisition
There is now a slit acquisition script available for NOTCam to speed up NIR spectroscopy. The script notcam.acquisition can be used for both WF-cam (R~2500) and HR-cam (R~5500) spectroscopy and is similar to the alfosc.acquisition script, but with a few additions. Due to flexure, an image of the slit is taken at every pointing and the position of the slit measured every time. Also, sky-subtraction of the acquisition image is available, if needed, when the source is fainter than the background.
2014-09-23 Helium hose leakage
Yet another helium leak in the old hose of the PTR cooling system. The hose was exchanged with the spare and the broken hose welded, tested, and pressurized. The location of the leak was again close to the instrument end.
2014-09-22 Temperature control board changed
A malfunctioning temperature control board in NOTCam was exchanged with the spare. The broken board will be repaired.
2014-09-04 Rotator limit check in scripts
The NOTCam template scripts notcam.9point, notcam.5point, notcam.beamswitch, and notcam.abba (see NOTCam scripts) have been upgraded to include a check on the time to reach rotator limit. We use maximum overheads estimate to calculate how long the scripts will take to execute. If this time is shorter than the time for the rotator to reach limit, the scripts will abort with an error message "[ERROR]: Rotator will reach limit before script ends. Aborts!" . The user should modify script parameters to lower execution time or turn the rotator by 360 degrees.
2014-08-28 New PTR hose ordered
We have ordered another 12.5 m long 20 mm diameter stainless steel braid over flexible (ultra-flex) helium hose from Kelvin International. It will be delivered with 99.9995% (UHP) Helium.
2014-08-19 Another helium hose leakage
Another helium leak in the old hose of the PTR cooling system. The hose was exchanged with the other old spare. The broken hose fixed, tested, and pressurized. The location of the leak was again close to the instrument end. The old hoses apparently break more frequently.
2014-04-14 PTR helium hose leakage
A major helium leakage in the hoses of the PTR cooling system was noted 9/4 and verified 11/4. On 14/4 the location of the leak was found to be in the old hose (the one used for the return flow), and it was then exchanged with the spare hose (the other old hose). The broken hose will be repaired and kept as a spare. NOTCam was kept cold all the time, though on LN2 only over the few days before the bad helium hose was exchanged with the spare.
2014-02-17 Internal focus with wavelength
Measurements of fwhm offsets at both wavelength extremes for the potential use of one fixed internal focus setting, shows that this is fully doable. A fixed setting would be needed for low resolution spectroscopy modes covering the whole range from 0.8 to 2.4 microns.
2013-11-04 New PTR hose installed
The new PTR hose was connected as a replacement for the most repaired hose (supply hose). The new hose is more flexible than the old ones.
2013-10-24 NOTCam Opening
21 Oct - 12 Nov 2013
NOTCam was set to warm up on October 21st. Detector temperature reached +7 deg after 3 days. Transported to the service building clean room on October 24th. Operations: cleaning pressure sensor, exchanging fill/vent tube o-rings, wet cleaning of entrance window, installing WF-camera imaging mask in the aperture wheel (taking out the 64 micron slit), installing replacement KG2 filter for out-of-spec KG4. NOTCam was reassembled and closed on October 28th and pumping started. Transported to the dome on the 31st, pumping continues. After one week of pumping the pressure is 2E-04 mbar at a temperature of 15 degrees. PTR switched on on Nov 8th, LN2 filled on Nov 11th.
2013-10-10 Illumination correction A section describing the WF-camera illumination correction has been added to the NOTCam Calibration page . There is a model available for the H-band, kindly made by Janz, Laurikainen, Lisker et al. 2013 who obtained the needed mapping of the field-of-view for the H-band in photometric hours during their own observing time. Mapping of the J and Ks bands is still pending.
2013-09-14 New calibration script notcam-calibs At the end of an observing night, when the dome is closed and dark, you can start the script notcam.notcam-calibs [date-string] to automatically take 10 darks of each of the used exposure modes during the night. See more info under NOTCam scripts.
2013-05-27 NOTCam warm-up and pump NOTCam was warmed up in the dome on May the 27th, being all the time on the pump. The LN2 tank was dried by blowing dry nitrogen gas trough it (gas from our LN2 flasks heated up with a bulb) and then pumped. NOTCam was cooled down again on June the 6th. We achieved a relatively good vacuum of NOTCam of 1E-05 mbar when starting to cool it with the PTR only. At a temperature of -80 degrees we started to fill with LN2 and the pumping was disconnected. The good vacuum was by this action lost to the more typical 2E-04 mbar, pointing again to thermal stresses at the fill tube O-rings. On the next opening these should be exchanged and we should increase the heating around these when filling. The behaviour of NOTCam is currently pretty much like it was before this exercise, but it is an important lesson that we actually can go for a warm-up without the need to open and bake the getter, a less risky procedure with less work.
2013-04-29 NOTCam reset-level monitoring The reset level (i.e. the ADU level of the pre-integration readout) is supposed to be a smooth function of detector temperature. At commissioning the DC-offset voltages were adjusted such that the level is 6000 ADU at operating temperature, a level that maximizes the dynamical range and minimizes bad pixels. Over time we have experienced jumps in this level, frequently so large that readjusting the DC-offset voltages is necessary. The first monitoring of the level (at every hour when NOTCam is off the telescope) started in June 2012. Because of the 1GB of data generated every day, the monitoring script was upgraded to analyse the images on-the-fly and delete them immediately. Monitoring re-started.
2013-04-20 NOTCam spare PTR hose arrived A 12m long 20mm diameter spare helium hose for the NOTCam PTR arrived from Kelvin International.
2013-02-01 Exposure time checking On rare occasions controller glitches have given exptimes shorter than requested. An "exptime-checker" program has been made to check every file automatically from now. An error message is displayed on the Talker messaging system if a problem is found.
2013-01-25 PTR - cooling water maintenance We had repeated problems over weeks in adjusting the PTR cooling water temperature. It was found that the PTR cooling water was dirty, so the whole secondary water circuit was properly cleaned. Also, the purg-o-matic valve has been exchanged for a new.
2013-01-22 Large reset level jump The dc-offset voltages had to be adjusted for the two lower quadrants, as the reset level had jumped from 6000 to 2000 adu. The reason for this is unknown.
2012-10-24 PTR hoses He leak fixed Again a Helium leak was detected in the PTR system. The leak originated in a crack in the flexible part of the supply hose, close to the instrument end, again near the previous point of repair, in a very vulnerable location where hoses are bent during operation. The tube was cut, re-welded, vacuum pumped and refilled with Helium.
2012-10-03 Upgrade of to v2.5 The IRAF quicklook reduction package has been upgraded to v2.5 in order to be able to handle both the old and the new fits header keywords. See download and installation info and NOTCam User's Guide - Data reductions for details.
Note that you need this new version to work on data taken after October 1st 2012!
2012-10-01 New fits header keywords Several NOTCam fitsheader keywords have changed name. This includes the old keywords APERTUR, FILT1, FILT1ID, FILT1POS, FILT2, FILT2ID, FILT2POS, STOP, GRISM, LENS, CAMERA that are now called NCAPRNM, NCFLTNM1, NCFLTID1, NCFLTPO1, NCFLTNM2, NCFLTID2, NCFLTPO2, NCSTPNM, NCGRNM, NCCAMNM, NCFOCUS . All related software has been updated and tested. Also, from now on we use editsetup to change the NOTCam instrument setup, and the old wheel definition files are obsolete.
2012-07-01 A new KG4 filter A new filter replacing the out-of-spec KG4 arrived from the provider at no cost. Installation is pending.
2012-06-20 New controller - User requirements The 34 page document User Requirements for the new Controller and Data Acquisition System for NOTCam (PDF file) was presented to the controller development group at CUO (Copenhagen University Observatory).
2012-06-14 Easy spectroscopy calibration The new script notcam.easy-calib is offered to make the in-situ spectroscopy calibration with NOTCam easier. Since this must be done at night and involves closing the mirror covers, the script easy-calib fid1 [fid2] [fid3] [fid4] [fid5] allows one to do 1-5 wavelength regions in one go. The script takes 1 Argon, 1 Xenon, and 3 Halogen lamp images per wavelength region (indicated with the FILTER ID number of the broad-band filters used as order sorters), each using the default exposure times. Calculate 2 minutes per region plus 1.5 minutes for mirror-cover closing and opening. The script uses notcam.setup-spec and notcam.calibexp. See more info under NOTCam scripts.
2012-04-20 Spectroscopic throughput in Z and Y Flux standards observed through the pinhole were used to derive the efficiency in spectroscopic mode of the new wavelength ranges available and defined by the broad band Z (0.830 - 0.927 microns) and Y (0.978 - 1.081 microns) filters. Adding up the two partly overlapping orders, the efficiencies are ~ 7% in Z and ~ 9% in Y, compared to ~ 10% in J. Remeasuring the single-order efficiencies of H and K gave 7% and 6%, respectively. To be verified with new observations.
2012-04-01 The BK7 and KG4 filters Calibrating the KG4 and BK7 glass filters that were installed in filter wheel 1 in order to be used in combination with the Z-filter to improve the red leak problem. BK7 reduces the thermal leak by merely 25%, while it also reduces the throughput in the Z-band by almost 10%. KG4 reduces the thermal leak by 90%, but it was found that its throughput in the Z-band was much lower than specification, and the provider was contacted.
2012-03-14 NOTCam Opening
14-15 March 2012 to install filters
A planned opening took place for the installation of filters: the two glass filters BK7 and KG4. The charcoal getter was baked for 24 hours, as usual. The opportunity was taken to bake the whole instrument at 37 degrees for 8 days. The LN2 tank was pumped as an attempt to diminish the suspected water content.
2012-03-04 PTR hoses He leak fixed A Helium leak was detected in the PTR system. The leak originated in a crack in the flexible part of the supply hose, close to the instrument end, near the previous point of repair, in a very vulnerable location where hoses are bent during operation. The tube was cut 10 cm, re-welded, vacuum pumped and refilled with Helium.
2012-01-17 Upgrade of to v2.4 The IRAF package for quick-look reductions of NOTCam imaging has been upgraded to v2.4. This version includes optional distortion correction in the tasks reduce and reduce_bs. A task mklincor for pixel-by-pixel correction of the non-linearity was added. The package is NOT a pipeline and was intended for quicklook reductions. See download and installation info and NOTCam User's Guide - Data reductions for details.
2012-01-12 WF-cam distortion model A model for the WF-cam optical distortion for the J, H and Ks filters based on high-quality data from 2009 is available in the form of IRAF transformation records to be used with geotran to correct the images. See NOTCam Calibrations for details. Optional correction is included in the NOTCam Quick-Look reduction package ( v2.4).
2011-11-30 Details on observing overheads Results from detailed investigations of the observing overheads can be found in NOTCam User's Guide - Overheads.
2011-10-12 Speeding up dithering For the expose command the telescope dither can now be performed as soon as the shutter has closed (using the new command notcam.wait-shutter-closed) i.e. the teloffset is done while reading out the array and storing the file, interlacing these overheads and increasing the speed of observing. This has been implemented for the all the NOTCam template scripts such as notcam.skyflat, notcam.5point, notcam.9point, and notcam.beamswitch, and it will be implemented in the NOTCam script generator. Timing measurements show that we now save on the average ~ 5.6 seconds per exposure in a dither sequence. The total time spent on a dither pattern is T_now/T_before = 0.76 for exp 3 and 0.82 for exp 10.
2011-07-01 PTR compressor oil filter replaced PTR maintenance: the compressor oil filter was exchanged with a new refurbished unit from Iwatani after a run time of 63199 - 28316 = 34883 hours.
2011-06-10 Z-band leak down The leak around the transparent edges of the Z-band filter was stopped by flipping the filter around in the holder and adding a cold stop ring to it. In total, there is now less background in the Z-band filter (the Z/Ks background ratio went from 1.2 to 0.88). As those numbers show, however, there is still a significant IR leak through the filter, and mainly beyond 2.5 microns, and we are currently asking quotes for a suitable blocking filter to mount in the other filter wheel. See updates in the ZY filter commissioning report.
2011-05-15 NOTCam Opening 24-31 May 2011 due to Filter-wheel-2 failure It was realized that filter wheel 2 started to fail on positions as described in Fault report 1514. No definite diagnosis could be made, and it was decided to warm up NOTCam for an opening (19/5). NOTCam was transported to the SB on the 24th and opened on the same day. It was found that the flexible coupling between the motor and the worm gear was broken. The reason for this remains unclear. Not only filterwheel 2, but all wheels therefore got their couplings exchanged with new ones. Also, all the gear wheel bearings were exchanged after 3.5 years use, as part of our maintenance plan. The Balzer pressure sensor was cleaned. We took the opportunity to flip around the Z-filter and put a cold stop on top of it. Instrument cleaned inside, charcoal baked, and NOTCam closed and set to pumping 31/5.
2011-03-18 NOTCam Cookbook (new format). As part of the new instrument specific NOT Cookbooks in general, the new format NOTCam Cookbook was added and thereby replaces the old NOTCam step-by-step manual for observers, which is now obsolete. Included in the new NOTCam Cookbook is also the preparing NOTCam observations page, a link which now becomes obsolete and will be removed. The NOTCam Cookbook printed version is found in an orange booklet in the control room.
2011-02-19 New Z and Y filters! The new broad band filters: Z (0.830 - 0.927 micron) and Y (0.978 - 1.081 micron) were commissioned (see report).
2011-02-12 New filter focus offsets. In conditions of good seeing, the opportunity was made after having had NOTCam open and disassembled, where special care was taken to align the detector plate when re-assembling, to re-measure the focus offsets between cameras and filters. The focusing procedure was found to converge quicker than it used to, and the scatter over the fov is smaller, probably owing to the better aligned detector and the re-calibration of the pyramid in late 2010. The new focus offsets are implemented in the instrument setup scripts notcam.setup-ima and notcam.setup-spec.
2011-02-12 Reduced overheads. Some internal overheads that had increased over the years have been taken care of. Both the Talker Log and the Obslog are now running on another computer, and we are back to nominal overheads for NOTCam data acquisition. See Overheads at the NOT and more details in the NOTCam User's Guide.
2011-01-11 NOTCam Opening 11-13 Jan 2011 NOTCam was opened on 11/1-11 for the installation of two newly acquired broad band filters: Z (0.830 - 0.927 micron) and Y (0.978 - 1.081 micron), see updated NOTCam filter list. The instrument was cleaned for dust, the 44 micron slit was aligned and fixed, a new PROM with modified code for the focus mechanism was installed and tested, the gear wheel bearings were checked, and the movement of filter wheel 1 was made smoother. The detector plate was well aligned with respect to the camera wheel housing when re-assembling the instrument. NOTCam was closed and put on the vacuum pump on the 13/1 with newly baked charcoal inside. A baking at <= 40 degrees of the whole cryostat was started thereafter. Functionality of detector and all wheels checked.
2010-11-18 QC analysis script upgraded The Detector Quality Control for NOTCam was upgraded. The analysis script was rewritten from IDL to python and some improvements and additions were made.
2010-10-28 Baffle calibration lamps online The NOTCam spectroscopy calibration unit, the three lamps (Halogen, Argon, and Xenon) mounted on the inside of the baffle lid, are now on-line and can be controlled from the TCS and the Sequencer. Their status goes to the NOTCam FITS headers. The script notcam.calibexp sets up the telescope for baffle lamp exposures, an action which involves switching the TV off (for protection) and closing the mirror covers, and takes the requested exposures with the selected lamp. Usage is documented in the NOTCam Cookbook.
2010-09-29 Pointing Scripts Use the new Pointing Script Generator to prepare scripts for NOTCam targets. This speeds up target acquisition and is particularly useful if you do larger than 60'' dither steps, which is typically the case for beam-switching, and/or if you wish to have the target oriented differently than the default for NOTCam. Both field orientation and guide area is taken care of by the script and you save time at the telescope.
2010-09-27 Non-linearity correction Pixel-by-pixel coefficient images are available for the correction of the inherent non-linearity of the NOTCam detector. The correction must be applied to the raw data as a first step in the data analysis. For more information see the page NOTCam Calibrations.
2010-08-20 Filters in the parallel beam A note was added to the NOTCam Filters page about the shift in wavelength of the transmission curve across the field of the detector for a filter mounted in the parallel beam.
2010-06-18 Leaking PTR hoses fixed The PTR hoses had suffered a longer time from a minor leak, implying refilling with Helium gas about once per month, but suddenly the problem deteriorated and needed repair. The leakage was located to be close to the instrument end, and the problem was fixed by cutting the hoses shorter. The NOTCam mounting instructions were modified to make sure the hoses are properly treated when mounting/dismounting NOTCam.
2010-05-27 Upgrade of to v2.3. The NOTCam quick-look reduction IRAF package on Florence has been upgraded to v2.3. The same for SLO computers. See the NOTCam User's Guide - Reduction section for documentation.
2010-05-25 Remotely controlled QC lamp The new remotely controlled lamp for NOTCam Quality Control, domeflats, and various tests has been commissioned and calibrated, and all documentation has been updated.
2010-05-09 One year cold! NOTCam has been kept at operating temperature (-199 degrees) for one year with no incidents.
2010-02-27 NOTCam instrument setup scripts. The instrument setup script for imaging mode notcam.setup-ima has been upgraded to include the focus offsets (i.e. tcs.focus-delta) corresponding to the filter and camera selected. NB! Note that the script assumes you have put the telescope focus for the WF camera and the Ks band as the tcs.focus-position value, which is also the default focus for NOTCam. There is a new instrument setup script for spectroscopy: notcam.setup-spec. Both scripts are described on the page NOTCam Sequencer Scripts.
2010-02-06 Focus pyramid re-calibrated. The focus pyramid for NOTCam, first installed and calibrated in 2005, was re-calibrated for both cameras with through-focus data taken in good seeing conditions in September 2009. Using the newly calibrated pyramid, the focus offset between the WF and the HR camera is now found to be +136 units (compared to earlier +90).
2010-01-14 New clock boards. The new clock-boards for the NOTCam controller were installed and tested. The old boards had clock voltage levels only ~3.5V while the new ones give 5V, according to specifications. This removed completely the shift register glow and provided much nicer images in low background mode (darks, spectra etc). See more info in NOTCam new clockboards. Also, the measured readnoise for the reset-read-read mode is both lower (8 electrons) and more stable, see the Detector Quality Control results.
2009-10-07 New LN2 filling nozzle The new LN2 filling nozzle, designed and made by Pedro out of a highly insulating special plastic material (same as used for the fill tube "cork"), was taken in use on the 20th of May. The vacuum five months after cool-down of NOTCam is on the average 1E-05 mbar (see newplot). One year ago using the old filling nozzle the average vacuum five months after cool-down was 5E-05 mbar (see oldplot).
2009-09-28 Improved guiding for large teloffsets New NOTCam guide areas for beam-switching were defined as NOTCAM-S-posX, NOTCAM-N-negX, NOTCAM-E-negY, and NOTCAM-W-posY, where the S, N, E, W sky directions for moving the telescope are valid only for default field rotation (-90), whereas the pos/neg XY directions for moving the star(!) on the array are always valid. Documented in the NOTCam Cookbook. The improvement from the earlier version from 2008 is basically larger areas and higher chances of finding a guide star, as well as easier use. The beam-switch observing script uses the same terminology.
2009-06-11 NOTCam remote observing NOTCam has been used in remote mode for the first time during the three nights 9/6 to 11/6 2009 by the Nordic-Baltic Optical/NIR and Radio Astronomy Summer School on "Star Formation in the Milky Way and Nearby Galaxies" held at Tuorla Observatory, Turku, Finland. The observations were successful, and all student groups got their about 2 hour observing slot with NOTCam each.
2009-05-28 Upgrade of some NOTCam SEQUENCER scripts The NOTCam sequencer observing scripts notcam.5point, notcam.9point, and notcam.beamswitch have been upgraded to higher flexibility. The user can select exposure mode (exp, mexp, frames), dither step size, an additional offset to allow for skewed grids (to avoid a fixed grid pattern and repeated positioning on the dead column), and there is an option to repeat the dither pattern a number of times. For the beam-switching mode the additional features are: selection of beam offset and direction on the sky.
2009-05-26 Upgrade of to v2.2. The NOTCam quick-look reduction IRAF package on Florence has been upgraded to V2.2. The reduction scripts reduce and reduce_bs now check the FITS header for the date of the observation in order to use the corresponding exact pixel scale for internal shifts as those applied by NOT2MEF and thus available in the FITS headers. The script dophot is made interactive and documented. See the NOTCam User's Guide - Reduction section for documentation.
2009-05-18 Pixel scale in FITS header The actually measured pixel scales for the two NOTCam cameras: 0.234"/pix for the WF camera and 0.078"/pix for the HR camera, were found to be slightly wrong in the FITS header keywords provided by the NOT2MEF program. The old values of 0.23"/pix and 0.08"/pix had been used up to now. NB! The correct pixel scales were always in use for slitoffset and teloffset calculations.
2009-04-10 NOTCam opening due to accidental warm-up An accidental warm-up after having been cold for 360 days triggered a complete thermal cycle of NOTCam. After baking the getter, closing and pumping, while testing the response of the instrument, it was discovered we had lost contact with quadrant #2. The pre-amplifier daughter board for this quadrant had broken, and eventually the fault was located to a transistor mounted inside the vessel on the detector PCB. New pre-amp boards were received from CUO on May the 3rd. The old one for quadrant #3 was moved to quadrant #2 and the new daughter board went in for quadrant #3. The array was taken out and put into a spare PCB. Based on warm tests we confirmed having contact with all quadrants. It turned out, however, that quadrant #3 had a funny behaviour realized only with cold tests. One of the resistors on the board (undocumented) had been given a wrong value. Replacing this one with the one on the original broken board solved the problem. NOTCam was ready for use on the 9th of May.
2009-03-08 Upgrade of to v2.1. The NOTCam quick-look reduction IRAF package on Florence has been upgraded to V2.1. The script now does the bad pixel correction internally and an extra option is made for suppression of the bias tilt (or "dc-gradient") caused by the reset anomaly. The scripts and will work even if there are no alignment stars available in the images, in these cases using the "sregister" task which is based on WCS header info, giving a slightly worse image registration. Trimming of images is set as an option. See the NOTCam User's Guide - Reduction section for how to use the package.
2009-01-19 NIR standard star catalogues NIR photometric standard star catalogues can now be directly loaded on the TCS. These are (Hunt et al. 1998, AJ 115, 2594) and (Leggett et al. 2006, MNRAS 373, 781). See NOT standard star catalogues and for more IR specific info see IR astronomy reference info.
2008-12-12 Differential twilight flats The stronger reset-anomaly introduced with the upgrade of the array electronics in Dec-2007 caused problems with the differential flats. Suppression of the dc-gradient (aka bias tilt) in the difference images was added as an option in the reduction software. The difference between using this correction option and not is shown in the NOTCam User's Guide - Reductions section on master flats.
2008-09-12 Sequencer Scripts and Documentation The NOTCam Sequencer Commands are documented in the NOTCam SEQUENCER Command Reference, and a set of simple observing scripts are documented in the NOTCam SEQUENCER Script Reference . These are available from any directory on the data acquisition computer.
As before, the user can also construct own scripts through the Observing Script Generator, storing the scripts in the user's directory on the data acquisition computer, (exceptions are beamswitch mode and scripts that involve changing filters, as well as more complicated scripts that can not be handled by the generator yet).
2008-06-09 Safe autoguiding for larger teloffset. In order to guarantee safe auto-guiding even for beam-switch observations, which improves both pointing and image quality, a set of smaller guide areas was defined for NOTCam. In the case of normal, small-step dithering (10-15''), the default area works fine, but if you teloffset larger steps, make sure you use the correct area. For more details, see the section on Autoguiding in the NOTCam Cookbook.
2008-05-22 Sequencer Mode The NOTCam Sequencer was offered in visitor mode for the first time on 22/5-2008. The NOTCam User's Guide and the NOTCam Cookbook for observers are both updated accordingly. The Observing Script Generator produces scripts in sequencer format.
2008-05-02 Holding time test NOTCam was left without filling for 14 days and the detector temperature was found to stabilize at -191 degrees (the ambient temperature being around 11 degrees). This indicates a very good vacuum, and is, in fact, the best we have obtained. In February 2007 after the improved vacuum behaviour upon cleaning the large vessel O-ring the detector stabilized at -185 degrees (the ambient temperature being around 7 deg).
2008-04-30 New Science Array images An RGB coded NOTCam image of the Ring Nebula (M57) was obtained by combining broad-band J and H images with narrowband H2 v=1-0 S(1). Beamswitch observations were performed and the data was reduced using the quick-look reduction tools in the IRAF package
2008-04-18 IMAGETYP keyword bug Most NOTCam images from 16/4 and 17/4, i.e. NCrd16* and NCrd17* data, are lacking a correct IMAGETYP keyword. Only the first image in a script has the IMAGETYP set correctly, and the remaining images have 'blank' entries. This was corrected 5/5-08, and the sequencer command for giving the IMAGETYP header (e.g. notcam.imtype FLAT) is again a sticky command.
2008-04-02 Entrance window crack Eventually, it was found that the entrance window had a tiny crack, and it was replaced by a spare entrance window. After this, the vacuum performance during warm (+20 deg) pumping of the cryostat was much better than ever experienced with NOTCam at the NOT, both on short and longer time-scales (reaching 6E-05 mbar after 2 weeks), see also pressure plot.
2008-03-10 NOTCam opening
Cryostat maintenance work
Trouble shooting and maintenance work on the cryostat was performed during Feb and March 2008. A leakage test found a problem close to the entrance window. The entrance window O-ring was therefore suspected and taken out to be cleaned. In addition, all other O-rings were cleaned: the large vessel O-ring, the fill/vent tube O-rings, the PTR cold head O-ring, the O-rings for all the connectors, etc. Also, the pressure sensor was cleaned. In addition, the whole cryostat was baked to a max temperature of 40 degrees - twice. First with the cold shield removed, in which case it was possible to reach only about 30 degrees. Then the cold shield was re-installed and the cryostat was baked at 37-40 degrees during 10 days. The vacuum reached was only 1.0E-02 mbar.
2008-02-12 New Science Array JHKs zeropoints The New Science Array sensitivity was measured on standard stars on the commissioning night 13-12-2007. We obtained the following zeropoints (for 1 e/s): J = 24.11, H = 24.06, and Ks = 23.45 mag, using fixed extinction coefficients in the fit. Note that the gain is 2.5 e/adu. The Exposure Time Calculator has been updated.
2008-01-17 IRAF package upgrade to v2.0 The IRAF package for quick look reductions, available on the Florence computer in the control room, has been upgraded to version 2.0. Any small step dither pattern can now be reduced with the script, and any beamswitch pattern with the script For a description see the NOTCam User's Guide.
2008-01-16 NOTCam opening
Vacuum problems
There is a problem with the vacuum of NOTCam, and we had to let it warm up by the end of December for a baking of the charcoal and an examination of the problem. No conclusive results. The next cool down was done while pumping all the time, both during the pre-cooling stage with the PTR and during the LN2 cooling. Disconnecting the pump was fatal, and it was decided to attach the pump to NOTCam on the telescope while observing for the coming observing runs in January.
2007-12-18 The PTR rotary valve is back! The PTR rotary valve is back from maintenance in Japan and was today mounted on NOTCam. Performed and cooled as expected.
2007-12-13 New Science Array commissioning (report) First cold tests with the new Science Array. Around 1 % bad pixels, but mainly in one corner. One bad feature. Read noise ~ 10 electrons, gain = 2.5 e/adu. Linear within 1% up to 20000 adu or more. Saturates at 56000 adu. Flatfield is much flatter than previous arrays. The general performance of the new array is described in the document The New NOTCam Science Grade Array (SWIR3).
2007-12-07 NOTCam opening
New Science Array installed!
The new Science Array was successfully installed in NOTCam. Warm tests verified that all quadrants worked well. Again the cool down was done directly with LN2 because the PTR was unavailable.
2007-12-03 NOTCam opening
Array electronics upgrade
The array electronics was upgraded. The bias voltages are set to recommended values and can also be adjusted. After many warm tests in the clean room, cold test were made with the engineering grade on the telescope. The detector quality control data shows that the new PCB works well in all aspects, but does not cure the dark current problem. NOTCam had been cooled down directly with LN2 since the PTR was sent to KIC for maintenance. Due to severe problems with the vacuum, no observations were made during night time. NOTCam was emptied of LN2 by putting the telescope to 6.4 degrees altitude and the rotator position to -65 degrees to prepare for another warm-up and opening needed to install the new Science Array.
2007-11-30 PTR rotary valve sent to KIC The PTR rotary valve was sent to KIC for maintenance. While it is gone we rely solely on LN2 for cooling.
2007-11-24 Exchange of bearings As a preventive means, all worm wheel bearings are exchanged with new ones every two years. This time all the old ones were found in a perfect state, but were still exchanged.
2007-09-30 PTR vibration tests PTR vibration tests on stars in 0.3'' seeing obtained with the NOTCam HR camera (0.078''/pix). The PTR was alternatively ON and OFF, and in each case a set of 5 consequtive and dithered 15 sec exposures of a star with high S/N was obtained. Autoguiding was on and the Ks band was used. There is no correlation between the measured FWHM of the stellar PSF and the status of the PTR.
2007-05-10 HR-camera optical quality The NOTCam HR camera delivers images of high optical quality. See the H-band combined image (N=9) of the central 60'' of The Globular Cluster NGC 4147 with FWHM = 0.44''. The ellipticity of stars is typically 0.02 to 0.05 all over the FOV. The pixel scale is found to be 0.078''/pix.
2007-04-01 Quick-look reduction package The IRAF package (v1.0) for quick-look reductions of NOTCam imaging is available on the florence computer in the control room. For a description check the NOTCam User's Guide.
2007-02-27 Second PTR rotary valve tested The second replacement rotary valve for the PTR sent from Kelvin International was mounted on NOTCam. It did not cool properly and had very strong vibrations. The old unit is back in place.
2007-02-13 The NOTCam cryostat performs well! Finally the NOTCam cryostat behaves well after a long time hunting for the micro-leak problems. The main cause of these problems was deformed o-rings in the fill and vent tubes because of temperature stresses when spilling LN2 at low altitudes as well as during LN2 filling. In addition, metallic dust had accumulated on the inside of the main vessel o-ring, realized only by taking the o-ring out of its rail. These problems have been fixed. NOTCam is now cooled only by the PTR during storage, and it stabilizes at a detector temperature around -183 deg without any loss of vacuum.
2007-01-17 NOTCam opening
The large vessel o-ring cleaned.
In November 2006 the vacuum was found to be too poor to keep going without another thermal cycle and baking of the getter before the next runs starting in the end of January 2007. NOTCam was warmed up and moved to the SB clean room 4/12. During this opening the opportunity was taken to work on the wheel controls. The o-ring was taken out and cleaned properly before putting it back in position with some vacuum grease. The getter was baked as usual and mounted hot just before closing and starting the vacuum pump. After 8 hours of pumping: 1.0E-02 mbar. After 24 hours: 2.4E-03 mbar. After 48 hours: 1.0E-03 mbar. After 3 days: 5.7E-04 mbar. (Dewar temperature still 20 degrees.) No vacuum loss during transport to the dome! After 4 days of pumping: 2.4E-04 mbar (no cooling). After 5 days: 1.3E-04 mbar. PTR started after 5 days and 4 hours of pumping.
2006-11-07 PTR rotary valve replacement unit The rotary valve replacement unit for the PTR arrived from Iwatani (Japan) and was tested on the instrument. Unfortunately, it failed to cool and had a strange out-of-phase sound (compared to the old working unit). The old unit is still working like a clock, but it has been running for more than 27000 hours while a replacement is recommended after 15000 hours.
2006-10-05 New Science Array arrived! The new Science Array from Rockwell arrived and is stored safely until installation in NOTCam.
2006-08-28 Improved filter focus offsets In good seeing conditions (FWHM 0.3''-0.5'') the focus offsets for the filters YnJHKs were determined for both the WF and the HR camera. These values can be found in the NOTCam User's Guide. The focus is wavelength dependent so these values should be good also for the narrow band filters.
2006-08-28 NOTCam BIAS upgraded Upgrade of the data acquistion software BIAS with the inclusion of a new clean-array command clear and some bug fixes.
2006-08-22 PTR compressor oil filter replaced PTR maintenance: the compressor oil filter was exchanged with a new refurbished unit from Iwatani after a run time of 28316 hours.
2006-07-12 NOTCam opening
Bad vacuum
Because of bad vacuum NOTCam was warmed up on July 12th and moved to the SB. Opened on 18/8. The inside of the vessel was cleaned. Charcoal baked. The procedure of mounting the getter was made easier in order to speed up the time from taking the getter hot out of the owen to the start of vacuum pumping. Closed on 22/8.
2006-06-12 Error in FITS keyword FILT2ID for J filter Note an error in the FITS header keyword FILT2ID for the J-band filter in the period January to June 2006, where the FILT2ID value, which refers to the NOT filter number, was 210 instead of the correct value 201.
2006-05-11 Light reduction Four masks are installed in the stop wheel for light reduction purposes. If you have a very bright target you can use a 9mm, 7mm, 6mm, or 5mm opening rather than the default 15mm one to get a throughput of 38%, 22%, 15%, or 10%, respectively. To be further tested.
2006-05-11 Engingeering grade array re-installed The engineering grade array was installed again in NOTCam on May 2nd. Its behaviour seemed normal in the warm state. Reached operating temperature on May 11th. Quality Control shows that gain is ~3 e/adu and readout noise ~10-12e, i.e. as before. Alignment OK. Updated zero-point magnitudes.
2006-04-22 NOTCam BIAS documented The NOTCam data acquisition software has been documented.
2006-april NOTCam opening
Science Array not well behaving
NOTCam was warmed up after the run 7-11 April and moved to the clean room and opened. A number of tests were made but nothing obvious was found. In the warm state none of the quadrants worked well, however. Cooling down again showed that all quadrants were bad even at operating temperature. The cause is unknown but the Science Array is useless and will be taken out.
2006-04-07 One quadrant dead Quadrant #3 was found non-responding on the 7th of April. The remaining three quadrants worked well and gave the normal values of read noise and gain. The observing run was modified to dither the targets within the three working quadrants. The reason for the problem is not understood. The instrument had not been opened since December and the new array worked well since installation. After the previous NOTCam run on March 15th there had been a case of a non intended warm-up because the PTR had turned off and the email warning monitoring of NOTCam temperatures had failed. NOTCam was pumped and cooled again.
2006-01-14 Science Array zeropoints The zeropoint magnitudes for NOTCam with the Science Array were found on the photometric night 14/1-2006 to be: 24.02, 23.97, and 23.33 for J, H, and Ks, respectively. Note that the zeropoints are for 1 e/s, not 1 ADU/s. The gain is 2.2 e/ADU. The Exposure Time Calculator was updated in April.
2006-01-14 First stars with Science Array Weather allowed for the first technical night on the sky with the Science Grade Array since its installation. A JHKs image of a stellar cluster can be seen here. The image scale was measured to be 0.234''/pix.
2006-01-14 Multi Extension Fits (MEF) format Commissioning of the MEF data format, the new User Interface software, and the upgrade of the data aquisition computer from elizabeth to marissa. The FITS keywords are upgraded. See New data formats and FITS headers.
2005-12-26 Vacuum losses remedied The origin of the problem of occational losses of vacuum upon LN2 spilling (at low altitudes and when the fill tube is lower than the vent tube) as well as during LN2 filling has been found. The vent tube extension is renewed with a thinner tube which is not in contact with the outer tube. The way the fill tube is blocked while observing is refined in order to avoid the inner tube getting contact with the outer tube, which causes temperature stresses on the o-rings and thereby vacuum losses.
2005-12-22 NOTCam opening
Vacuum problems
NOTCam was set to warm up on the 22/12 after an observing run of Venus at very low altitudes. It was found that thermal stresses on the vent tube o-rings caused vacuum leaks when LN2 was spilling out from the vessel as its orientation would be at very low altitudes.
2005-10-26 Science array commissioning (report) First cold tests of the Science array was made on the 25th of October. NOTCam was then mounted on the telescope on the 26th. Cosmetics, flat field, linearity and QE is much better than for the engineering grade array. For details check the NOTCam Science Array page. Bad weather did not allow opening the dome. Zeropoints and backgrounds must be checked later.
2005-10-20 Science array installed! The engineer grade array was taken out and the science array was installed in NOTCam. The procedure has been documented with pictures here (limited access). Tests performed to verify that the array can be read out.
2005-Oct NOTCam opening
Exchange of bearings
+ other work
All worm wheel bearings were exchanged. The new bearings were first lubricated with MOS2 - this time while rotating the bearings to get a homogeneous layer. All wheels tested and worked fine. The entrance wheel did not turn smoothly, and wheel teeth were marked by the strong pressure from the spring between the two halves of the worm gear. This was fixed by modifying some spacers. The Balzer pressure sensor was cleaned. The LN2 fill and vent tubes were opened for the first time and examined. Some oxidized parts were cleaned. The outer o-ring in each tube was found deformed. This could be the cause of some recent vacuum losses occuring every time we fill with LN2. New o-rings were mounted. The 4 narrow-band Östlin filters were taken out.
2005-07-20 NOTCam opening
Poor vacuum
Because of poor vacuum and holding time we decided to warm-up NOTCam. The holding time became merely 2 days and kept decreasing. NOTCam was pumped on the 9/8 while at a temperature of 15 deg C and a starting pressure of 37 mbar. After one hour 2.5e-03 mbar, and after 24 h 3.2e-04 mbar was reached. The PTR was started after 51 h of pumping. 12 h later the array was at -22 deg C in a vacuum of 4.78e-05. LN2 was filled 9 h after this, when the array was at about -50 deg C, and it took 24 hours more for the array to reach -196.3 (vacuum 7.6e-05 mbar). The holding time could not be properly checked, but was at least better than 3 days.
2005-Jun No increase of dead pixels The previously mentioned large number of "dead pixels" (that in addition seemed to come and go) were not actually dead! Investigation of the reset frame (i.e. the pre-read) showed that its peak histogram level had drifted from about 4000 ADUs towards zero. Pixels with zero or negative values in the reset frame, would end up as zero valued pixels in the resulting image (i.e. the end-of-integration frame minus the reset frame) because the program would set them so. This caused an apparent presence of dead pixels. Only slight modifications to the bias offset voltages were needed to adjust the reset frame level back to its original state. Thus, the bottom line is that the actual number of dead pixels in this array is around 1%, and that this number has been all the time the same. We suspect the origin of the problem is some drift in the array itself. And from now on we will regularly monitor the reset frames.
2004-Nov NOTCam opening
Focus pyramid installed
NOTCam has now a focus pyramid, similar to the one in ALFOSC, installed in the grism wheel. It is made of IR grade fused silica (Heraeus Infrasil 302), cemented together with a vacuum grade cement and held together by a spring metal frame. The pyramid post-processing software calculates the offset in telescope focus units from the correct focus and needs in principle only one image obtained. Makes NOTCam observing a lot more efficient.
2004-02-12 No further dead pixels When NOTCam had reached operating temperature on 12/2 a number of darks were taken to estimate the number of dead pixels. There has been a continuous increase with every thermal cycle during 2003. To our positive surprise there was no increase between now and the previous opening. The number of dead pixels measured on the 12/2-2004 was on the contrary only 2.5%. Dead pixels seem to come and go.
2004-02-05 Closing and leak test The anodized masks made to limit scattered light for both HR and WF imaging mode were installed. The 128 micron WF slit was taken out. Screws and nuts put on the 64 micron WF slit, but not strongly tightened. Then all wheels tested to initialize well, and all temperature sensors checked. The pressure sensor was replaced with a spare (old Hirac). The hot getter installed and NOTCam closed and set to vacuum pump. He leak test performed: no leak (i.e. value less than 2.0E-8 units on RGA). Continue pumping.
2004-01-29 NOTCam opening
Leak test
We made a leak test of NOTCam with He gas and the Residual Gas Analyzer. No leak was found, and we concluded that the pressure sensor is unreliable and will be replaced by a new one as soon as possible. (From now on we will perform leakage tests upon every closing of NOTCam to make sure the vacuum will be OK before transporting it to the dome. A leak testing procedure (limited access) has been put on the web. Then NOTCam was opened and the three new Ostlin filters were mounted. The Grism2 filter was taken out.
2003-12-10 Dead pixel evolution NOTCam has been warmed up six times during 2003, and the number of dead pixels has increased from 1% in Jan-2003 to 5% in Dec-2003 (given as percentage of dead pixels of the total 1024x1024 array). The detector is still the engineering grade array. The dead pixels mainly cluster along the left edge.
2003-11-10 NOTCam opening
Exchange of bearings
NOTCam was opened in the clean room for the very first time, as it now has its own table with a lifter such that the lid can be hooked up from the ceiling. We have exchanged all the worm wheel gear bearings to avoid further hick-ups. The new bearings have been lubricated by MOS2. Some of the old bearings were found to be very worn. A document was made on How to exhange bearings (limited access). Two recently arrived filters (Z and He IA) were installed. The focus mechanism was taken apart to find out what type of bearings are used.
2003-09-05 Pick-up noise removed! Over the summer months the work on reducing the persistant pick-up noise gave results. The readout noise is now close to specs for both readout modes, and the stripes are almost completely gone.
2003-08-19 NOTCam opening
Timeouts on filterwheel 1
Problems with stuck filterwheel during the service observations forced another warm up and opening. One bearing in the worm wheel drive of filterwheel 1 was this time exchanged. 10 days is about what is needed for a full warm-up, opening and cool-down cycle for NOTCam. See Fault report (limited access).
2003-08-07 NOTCam spectroscopy: first light! First light for spectrosopic mode! Check Commissioning report for status and progress.
2003-06-20 NOTCam opening
Grism #1 installed
The first NOTCam grism was installed in the grism wheel, and a grism holder mask was made. The Ostlin filters were taken out and found to have cracked top coatings. Returned to the provider. NOTCam closed 24/6, transported to the dome 26/6, PTR turned on 30/6, filled with LN2 1/7, and cold by 3/7.
2003-05-12 NOTCam opening
Installing Östlin filters
Filters arrived in the last minute and got installed, instrument closed 12/5. Problems in the startup of the pumping procedure, and the instrument barely reached operating temperature when observations started on the 17/5. The cold table temp. sensor had been broken.
2003-04-22 Stable arrray temperature Since using the PTR also with NOTCam mounted at the telescope, the array temperature is now much more stable throughout the night. It has been measured to deviate only 0.1 degree over the whole night for typical nights. This of course depends on the amount of LN2 that might be spilled.
2003-04-07 NOTCam opening
Filterwheel 2 + other work
The problem with filter wheel 2 was located to one of the wheel bearings, which had become stuck. A spare one was cleaned and then lubricated with MOS2 to replace the broken one, and the wheel now moves fine. The Balzer pressure sensor was cleaned according to this method (limited access only). One filter in filterwheel 1 which had become a bit lose in its holder was tightened by putting in a new copper spring. NOTCam was closed and started pumping on 2/4 after having verified that all wheels initialized well. The getter was baked again at the JKT at 150 degrees for about 4 hours. It takes 75 minutes before the pump reaches operating speed at 75 KRPM. By then the pressure was down to 1.2 10-1 mbar. After 4 hours pumping, P = 5.1 10-1 mbar. Transported to the dome on 4/4. Pumping continued. PTR started on 5/4. LN2 filled morning 6/4. NOTCam reached operating temperature 7/4 around midday.
2003-03-27 Warm-up time ~ 3 days Verified that the init switches worked OK and concluded that the sounds from the wheel and motor indicated a mechanical problem. Decided to open NOTCam. Found that by moving NOTCam early to the SB, warm-up time is about 3 days from turning off the PTR. The last day of LN2 filling was 5 days before turning off the PTR.
2003-03-17 Filter wheel timeouts Timeouts on filterwheel 2 during a visitor run. Could be initialized only using tty to run the UIF and on repeated trials. Was run with a fixed filter setup for each night the rest of the run. See Fault report (limited access only).
2003-03-12 PTR vibration tests No technical problems. New PTR vibration test data obtained, and no deterioration found down to a best seeing of 0.5'' FWHM. Student training. New linearity test data obtained for both readout modes. Photometric standards in broad and narrow band filters. One night lost to useless weather. Testing of the "buzzer", the device which is designed to monitor the shutter behaviour and give warning alarms if it fails.
2003-02-24 Leak testing using RGA NOTCam was transported back to the dome and tested for leakage using Helium gas and the RGA. No leaks found, except for a very minor leak around the amplifier boards, which has probably always been there.
2003-02-20 to 21 NOTCam opening
Slits and masks installed
NOTCam was opened. Getter baked in owen at JKT. Slit mounts designed and fabricated. Two slits mounted. Four hartmann masks mounted. Dummy grism test mounted. Two redundant filters taken out and two new filters installed. Collimator position marked. Pressure sensor found to be unreliable, a factor of 10 wrong. Should be cleaned next time. The roughing pump in the new pump is not capable of pumping NOTCam.
2003-01-16 PTR long hoses installed PTR long hoses installed. First test with PTR running when NOTCam is mounted on the telescope. Vibration tests showed no effect of the PTR down to a best seeing of 0.9'' FWHM for this night. Photometric standards. Snow and ice on the hatch impeded opening upper hatch all the way, and zenith distances less than 15 deg had to be avoided. No problems with the long hoses at any rotator/altitude positions.
2002-09-22 to 25 Commissioning Four half nights of which 20% was lost to bad weather. Focus tests of both cameras. Standards for both cameras. Experienced once that the shutter was hanging. Some narrowband filter testing.
2002-09-16 11th commissioning run (1/2n) Photometric weather. Photometric standards observed in JHKKs with both cameras, also across the FOV to check the photometric scatter. Check results.
2002-August Warm-up and pump NOTCam was brought to the service building and warmed up. No opening became necessary. Was pumped and cooled again.
2002-06-01 10th commissioning run (1n) First photometric night since shutter was OK! Many hours lost to miscellaneous problems as this was the first night after the M1 and M2 aluminisation procedure. Pointing model observations were made, as well as StanCam zeropoints. One photometric standard observed in H and K with the WF camera.
2002-05-13 Six months cold NOTCam has been kept cold for half a year and it is clear that during storage the detector is kept stable at -183 to -190 degrees by the PTR alone. Variations follow the variation in ambient temperature.
2002-04-01 First Visitor run! NOTCam was offered in visitor mode for the first time. The observing run was 7 nights. No time lost to problems! Half of the time lost to rain and snow, however. Juke box automatically burning around 26 CDs per night.
2002-03-29 to 31 9th commissioning run (3n) Totally lost to rain. The CD burning juke box was ready for use.
2002-03-18 NOTCam BIAS BIAS bug fixes and changes.
2002-01-26 to 28 8th commissioning run (3n) Mostly lost to thick clouds, some data taken through thinner clouds.
2002-01-02 to 03 7th commissioning run (2n) First night useless weather. Second night commissioning test data obtained. No technical problems.
2001-12-25 Shutter now fixed!
6th commissioning run (5n)
Useless weather. Shutter now finally behaves perfectly. No timeout on any wheels. First reliable noise and linearity test data obtained!
2001-11-30 Shutter speed modified The time it takes to rotate the shutter 90 degrees was changed from 100ms to 150ms. Extensive testing shows this seems to have solved the problem. A closed loop solution is being looked into where the shutter position verification is done automatically.
2001-11-27 5th commissioning run (2n) Shutter more unreliable than ever. Observations were repeatedly interrupted during the first night. The second night NOTCam was run with the shutter disabled. Some commissioning proposal targets done.
2001-10-29 NOTCam opening The stuck filterwheel was caused by the nut of a screw that had become stuck in the motor gear. All wheels, shutter and focus run well during extensive testings before closing NOTCam.
2001-10-07 4th commissioning run (2n) No data obtained. Useless weather. Stuck filterwheel. Shutter does not operate well.
2001-09-08 3rd commissioning run (4n) The high resolution camera was installed for this run and the two cameras are found to have foci almost within the range of instrument focus travel at the same telescope focus. A small modification will be necessary to ensure that it is not necessary to change telescope focus when changing cameras. All filters are installed. This run made it clear that normal operations will require two observers to make optimal use of the instrument and the NOT archive system need vastly increased bandwidth (a project which is already underway). More commissioning-time science programmes were executed.
2001-08-20 NOTCam opening
Filter wheel stuck
HR camera installed
Installation of the HR camera.
2001-08-10 2nd commissioning run (2n) The instrument was kept cold between the first two commissioning runs to allow us to gain experience with long term storage. A very small leak appears to be present. During this period, a number of software problems were resolved. The second commissioning run itself was less successful than the first with a stuck filter wheel and loss of synchronisation of the shutter.
2001-07-01 1st commissioning run (2n) The first commissioning run was very successful with no critical failures of the instrument at all. A considerable amount of verification data was collected and we were even able to collect some scientifically useful data. The biggest problem identified was how to prevent liquid nitrogen spilling from the dewar. Some small modifications may be necessary.
2001-06-30: FIRST LIGHT! 30 days of intense effort paid off with the first light image being on target and in focus, greeted with applause in the control room and broad smiles all around.
2001-06-30: NOTCam is mounted on the telescope It fits!
2001-06-15 First "recognisable" images on La Palma Give a geek a camera costing several hundred thousand Euros, and what will he do with it...?
2001-06-01: NOTCam arrives on La Palma NOTCam arrives on La Palma, apparently undamaged, and work begins immediately to bring the instrument into condition for use at the telescope.
2001-03-30: First recognisable images The new engineering grade detector has been installed in Copenhagen and the dewar cooled without a hitch. The detector was then used to obtain recognisable images through both cameras, the first such images obtained with this instrument. Final electromechanical and cryogenic tests are underway; the next step is to align the optics.
2001-03-19: Status NOTCam is still in Copenhagen. The mechnical destruction of the engineering grade detector late last year forced a new delay while we redesign the detector mount to avoid breaking the new one as well. The precise cause of the breakage is not known, so we are being careful to make sure that all possibilities are covered. We cannot, at this time, predict when NOTCam will arrive in La Palma nor when it will be available to the community, but we are doing our best to proceed as quickly, and as safely, as possible.
2000-03-09 Electromechanics, PTR & Controller All electromechanical systems now function correctly.
Leaks have been found in the Iwatani PTR hoses and are being repaired in the USA.
Controller work continues with highest priority at CUO.
2000-02-28 SPIE paper Paper submitted to SPIE describing NOTCam and current progress: "SWIR at the Nordic Optical Telescope: NOTCam", Abbott et al. 2000, SPIE 4008, p. 714-719 . A poster will be presented at the SPIE meeting in Munich at the end of March.
1999-11-15 Controller & MUX The MUX has been run in the test dewar, producing an image.
1999-11-01 3rd Cooldown Detector mount now achieves operating temperature, but very slowly (72 hrs). Pressure in dewar still increases while closed (stable at ~10-5 mbar). No contamination on window. No heat path through window mount. Cooling power is ~15W. LN2-only hold time: ~44hr.
Before cooldown: PTR wick installed, anodization cleaned off PTR & window flanges and cold shield, cooling braid to detector mount installed, Nextel removed from hemispherical baffle cover, dewar not baked(?).
99-10-15 Warmup Warmup rate is 38°/day. 6 days from cold to ambient. Contaminant haze present on window. No contamination on witness slides.
99-10-11 2nd Cooldown Detector mount does not achieve operating temperature, but sensor later found to be compromised. Four 25l bottles of LN2 used for initial cooldown. Cooling power: ~22W. LN2 hold time: ~31hr. Some cooling of window flange noted. Fill tube found to be potentially dangerous. RGA spectrum while cold appears clean.
Before cooldown: charcoal getter installed, hemispherical baffle trimmed by ~3 mm, dewar not baked.
99-09-30 1st Cooldown Lowest uncooled pressure: ~5×10-5 mbar. Condensation on window. Motors run while cold.
Before cooldown: standard 2-day bake.
99-09-29 CUO/NOTSA contract addendum mailed out
Back to top Last modified: April 19 2024